# Is GW190521 a gravitational wave echo of wormhole remnant from another universe?

### v1

@@ Abstract
A particularly compelling aspect of the GW190521 event detected by the LIGO-Virgo collaboration is that it has an extremely short duration, and lacks a clearly identifiable inspiral phase usually observed in the binary black holes (BBHs) coalescence. In this work, we hypothesize that GW190521 might represent a single, isolated gravitational wave (GW) echo pulse from the wormhole, which is the postmerger remnant of BBHs in another universe and connected to our universe through a throat. The ringdown signal after BBHs merged in another universe can pass through the throat of wormhole and be detected in our universe as a short-duration echo pulse. Our analysis results indicate that our model yields a network signal-to-noise ratio comparable to that of the standard BBHs merger model reported by the LIGO-Virgo collaboration. Though the Bayesian factor slightly prefers the standard BBHs merger model, it is not significant enough to rule out the possibility that the echo-for-wormhole model is a viable hypothesis for the GW190521 event.
I Introduction
@@
As of now, the LIGO-Virgo collaboration has detected nearly one hundred gravitational wave (GW) events Nitz:2021zwj ; KAGRA:2021vkt . Their successful detection has opened new avenues for comprehending our universe, particularly physics in strong gravitational fields.
@@
To date, most confirmed GW events have been confidently attributed to the mergers of binary black holes (BBHs) or binary neutron stars (BNSs), with their signals displaying the well-understood inspiral-merger-ringdown (IMR) morphology Nitz:2021zwj ; KAGRA:2021vkt . However, among the cataloged events, GW190521 is an unusual and intriguing case that might challenge this standard paradigm LIGOScientific:2020iuh , since it has an extremely short duration (the order of 0.1 seconds) and lacks a clearly identifiable inspiral phase usually observed in the mergers of BBHs. This lack raises questions about its physical origin. Though the LIGO-Virgo collaboration has interpreted GW190521 as the merger of BBHs with masses in $85^{+21}_{-14}M_{\odot}$ and $66^{+17}_{-18}M_{\odot}$, resulting in a postmerger remnant of $142^{+28}_{-16}M_{\odot}$ LIGOScientific:2020iuh , see also relevant e.g.Refs. Fishbach:2020qag ; Nitz:2020mga ; Estelles:2021jnz ; Capano:2022zqm , it is in tension with established astrophysical models of stellar evolution. Thus some alternative interpretations involving such as primordial black holes, cosmic strings, new light particle, and even horizonless compact objects have been presented Clesse:2020ghq ; Aurrekoetxea:2023vtp ; CalderonBustillo:2020fyi ; Sakstein:2020axg ; Abedi:2021tti ; CalderonBustillo:2022cja .
@@
In parallel, to investigate possible macroscopic signatures of quantum gravity and the longstanding black hole information paradox, a variety of horizonless exotic compact objects have been investigated, e.g.Cardoso:2019rvt for a recent review. The wormhole represents such an object connecting either two separate universes or two distant regions in a single universe through a throat Morris:1988cz ; Morris:1988tu , see also Poisson:1995sv ; Hochberg:1998ha ; Armendariz-Picon:2002gjc ; Garcia:2011aa ; Kanti:2011jz ; Damour:2007ap . It has been shown that if the remnant object after BBHs merged is a wormhole, the corresponding ringdown burst propagating toward the other side of wormhole can be reflected back by the photon sphere on the far end, yielding a sequence of delayed pulses, known as echoes Cardoso:2016rao . The properties of echoes, such as their amplitude and time delay, carry information about the underlying physical characteristics of corresponding wormholes, e.g.Cardoso:2017cqb ; Bueno:2017hyj ; Wang:2018mlp ; Correia:2018apm ; GalvezGhersi:2019lag ; Li:2019kwa ; Bronnikov:2019sbx ; Liu:2020qia ; Yang:2024prm .
@@
II Model
II.1 Our echo-for-wormhole model
@@
As a proof-of-principle model, we consider a Schwarzschild-like Morris-Thorne wormhole Morris:1988cz ; Morris:1988tu , in which the Schwarzschild metrics of both sides are glued at $r=r_{0}>r_{S}$, and $r_{0}$ is the radius of wormhole throat and $r_{S}$ is the Schwarzschild radius. The relevant issue on post-ringdown GW signals can be regarded equivalently as a one-dimensional scattering problem with a frequency-domain wave equation:
$\frac{d^{2}{\tilde{\Psi}}}{d{r_{*}}^{2}}+\left(\omega^{2}+V(r_{*},\omega)\right){\tilde{\Psi}}=0$ (1)
@@
In our model, see Fig.1, the BBHs inspiraled and merged in the left-side universe, and their postmerger object is a Schwarzschild-like wormhole connecting it to our universe, as a result the postmerger GW ringdown signal not only propagates outward at the speed of light, but also enters the wormhole throat and penetrate the right-side barrier into our universe. It can be expected that the reflections of such a pair of mirror barriers for the ringdown pulse will yield a series of echo pulses with sequentially suppressed amplitudes. The corresponding signals consist of a sum of wormhole quasinormal modes Bueno:2017hyj
$\Psi(t)\approx e^{-i\omega_{0}t}\left(\sum_{n=-\infty}^{+\infty}c_{n}e^{-n\pi t/L}e^{{\rm Im}(\omega_{n})t}\right),$ (2)
@@
$L\simeq 2\int_{r_{0}}^{\frac{3r_{S}}{2}}\frac{dr}{1-r_{S}/r}$ (3)
represents the separation between the mirror barriers, $\omega_{n}$ is the quasinormal frequency (QNF) of wormhole, and the coefficients $c_{n}$ can be written as
$c_{n}={\frac{1}{2L}}\int_{0}^{2L}dt\Psi_{1^{\rm st}{\rm echo}}e^{i\omega_{n}t},$ (4)
where $\Psi_{1^{\rm st}{\rm echo}}$ is the waveform of the first echo. The interval between two echoes is approximately $\Delta t_{\pm echo}\simeq 2L$, however, if the postmerger object is a wormhole collapsing into a black hole, the change of interval after a period is
$\Delta t_{\rm echo}\simeq{2L\over 1-{\Delta L\over L}},$ (5)
@@
The postmerger ringdown signal closely resembles that of a black hole with Schwarzschild radius $r_{S}$. As argued in Ref. Bueno:2017hyj , the waveform of the first echo is primarily governed by the wormhole quasinormal frequency (QNF) that lies closest to the corresponding black hole QNF $\omega_{0}^{\rm BH}$, thus it is natural to model the first echo as a Gaussian-like wave packet
$\Psi_{1^{\rm st}{\rm echo}}\sim e^{i\omega_{0}^{\rm BH}(t-t_{*})}e^{-{(t-t_{*})^{2}\over 2\beta^{2}}},$ (6)
where $\beta$ characterizes the width of the Gaussian-shaped echo pulse.
II.2 Template of waveform
It can be postulated that, after the merger of binary black holes, a postmerger wormhole forms as an intermediate state, which rapidly pinches off and eventually collapses into a black hole Wang:2018mlp , or that the amplitudes of subsequent echoes fall below the sensitivity threshold of current detectors Abedi:2021tti . In either case, only the first echo pulse can be detected in our universe [^3] Recently, numerous studies have focused on the search for echoes after the BBHs inspiraled and merged in one single universe, e.g.Abedi:2016hgu ; Westerweck:2017hus ; Lo:2018sep ; Nielsen:2018lkf ; Wang:2019szm ; Uchikata:2019frs ; Wang:2020ayy ; LIGOScientific:2020tif ; Ren:2021xbe ; Abedi:2021tti ; Abedi:2022bph ; Wu:2023wfv ; Uchikata:2023zcu , showing no evidence for echo, however, here what we consider is to take the GW190521 event itself as an echo (detected in our universe) of postmerger wormhole in another universe in which the BBHs inspiraled and merged.. According to (6), we use a simplified sine-Gaussian template:
$h(t)$ $\left(\sim e^{2\pi if_{c}(t-t_{c})}e^{-{(t-t_{c})^{2}\over 2\beta^{2}}}\right)$ (7)
$=$ $\mathcal{A}\mathrm{cos}\left(2\pi f_{c}(t-t_{c})+\varphi\right)e^{\frac{(t-t_{c})^{2}}{2\beta^{2}}},$
@@ II.2 Template of waveform
The amplitude $\mathcal{A}$ depends on the energy of the GWs emitted by the BBHs merger in another universe. However, in our case, the total energy emitted by the source is unknown, which causes the amplitude parameter inferred from the Bayesian analysis to encode both the intrinsic strength of the source and its distance to the detector. To facilitate the parameter estimation, we fix a reference strain amplitude $A_{ref}=10^{-19}$ at a fiducial distance of 1Mpc, the physical amplitude $\mathcal{A}$ is then rescaled by a dimensionless factor $A_{m}=\mathcal{A}/A_{ref}$ and set as a free parameter in our analysis. This normalization not only improves numerical stability but also yields a consistent scale for comparing with LIGO-Virgo BBHs models.
III Results
III.1 Our echo-for-wormhole model
@@
Table 1: Prior ranges of parameters for our echo-for-wormhole waveform (7). Here, $\psi$ is the polarization, while $\iota$ is the inclination refers to the orientation of binary relative to the line of sight. Geocent time parameter specifies the arrival time of the echo pulse at the Earth. The physical interpretations of other parameters are presented in Sec.II.2
@@
The priors of our echo-for-wormhole model are shown in the Tab.1. In the initial parameter estimation, we find that the maximum-posterior sky location inferred is consistent with that obtained using the LIGO-Virgo BBHs model, and also with the localization reported for ZTF19abanrhr, as shown in Appendix A. Thus we fix the location parameters with ZFT19abanrhr localization (RA, DEC)=(3.35847,0.60781)=(3.35847,~0.60781) to break the degeneracy between arrival time and sky position. The marginalized posterior probability distributions for all parameters in our model are presented in Fig. 2, where $f_{c}=56.93^{+1.63}_{-1.86}$ Hz and $\beta\approx 0.02$. This result is consistent with the narrow-band feature of an echo pulse produced by a wormhole in the frequency domain.
In Fig.3, we plot the bestfit echo-for-wormhole waveform (7) with the strain data in L1 detector in time-domain, and also the BBHs waveform for a comparison, where owing to the short duration of GW190521 (approximately 0.1 s), we additionally bandpass the strain data between 20-256 Hz to suppress high frequencies, results in an almost 5 cycle waveform. Here, to facilitate a direct comparison between the two models, we normalize both the strain data and the corresponding waveforms using the noise power spectral density $\tilde{S}_{n}$ provided by the LIGO-Virgo collaboration:
$\tilde{h}_{w}(f)=\frac{\tilde{h}(f)}{\sqrt{\tilde{S}_{n}(f)}},$ (8)
where $\tilde{h}_{w}(f)$ denotes the whitened waveform in the frequency domain.
@@
III.2 Bayesian comparison with BBHs model
Figure 4: The posterior distribution of intrinsic parameters and luminosity distance for the standard BBHs IMR event with IMRPhenomXPHM waveform. This distribution is consistent with the result reported by the LIGO-Virgo collaboration for GW190521 using the same waveform model Estelles:2021jnz .
Here, we use the IMRPhenomXPHM as the BBHs waveform template [^4] Though the waveform models NRSur7dq4 and SEOBNRv4PHM LIGOScientific:2020iuh might yield more precise parameter estimates, the differences between different BBHs models are negligible compared to the impact of distinct source hypotheses on the Bayesian analysis., as it accommodates high mass ratios and spin precession, effectively covering the parameter space relevant to GW190521. In addition, its frequency-domain construction offers high computational efficiency, facilitating a systematic comparison with our echo-for-wormhole model. We adopt the same prior distributions as those used in Ref. Estelles:2021jnz , this choice places greater emphasis on unequal–mass systems, thereby improving sampling convergence in the unequal–mass regime. And using the IMRPhenomXPHM waveform we recover posterior of relevant parameters [^5] Here, we also used the localization reported for ZTF19abanrhr, see Appendix A. However, we have verified that the results are actually nearly same with those without the ZTF19abanrhr localization. (see Fig.4), which consistent with the parameter estimates reported by the LIGO-Virgo collaboration Estelles:2021jnz , the merger of BBHs with masses of $m_{1}=97^{+32}_{-21}\mathrm{M}_{\odot}$ and $m_{2}=59^{+22}_{-25}\mathrm{M}_{\odot}$ at a luminosity distance $d_{L}=3.5^{+2.4}_{-2.0}$Gpc.
To evaluate the plausibility of our model as a potential explanation for the source of GW190521, we compute the SNRs of the BBHs model and our model for each detector, as well as the network SNR, by using the publicly available PyCBC software Usman:2015kfa . The results are summarized in Table 2. The network SNR is defined as the quadrature sum of the individual SNRs across the detectors,
$\text{SNR}^{2}_{\text{net}}=\sum_{ifo=\text{H1,L1,V1}}\text{SNR}^{2}_{ifo},$ (9)

### v2

@@ Abstract
A particularly compelling aspect of the GW190521 event detected by the LIGO–Virgo–KAGRA (LVK) collaboration is that it has an extremely short duration, and lacks a clearly identifiable inspiral phase usually observed in the binary black holes (BBHs) coalescence. In this work, we hypothesize that GW190521 might represent a single, isolated gravitational wave (GW) echo pulse from the wormhole, which is the postmerger remnant of BBHs in another universe and connected to our universe through a throat. The ringdown signal after BBHs merged in another universe can pass through the throat of wormhole and be detected in our universe as a short-duration echo pulse. Our analysis results indicate that our model yields a network signal-to-noise ratio comparable to that of the standard BBHs merger model reported by the LVK collaboration. For GW190521, Bayesian model selection yields $\ln\mathcal{B}^{\text{Echo}}_{\text{BBH}}\simeq-2.9$, indicating that the data favor the BBH hypothesis over our echo-for-wormhole model.
1 Introduction
@@
As of now, the LIGO–Virgo–KAGRA (LVK) collaboration has reported 218 gravitational wave (GW) events [2, 3, 4, 5, 6]. Their successful detection has opened new avenues for comprehending our universe, particularly physics in strong gravitational fields.
@@
To date, GW events have been confidently attributed to the mergers of binary black holes (BBHs) or binary neutron stars (BNSs), with their signals displaying the well-understood inspiral-merger-ringdown (IMR) morphology [2, 3, 4, 5, 6]. However, among the cataloged events, GW190521 is an unusual and intriguing case that might challenge this standard paradigm [7], since it has an extremely short duration (the order of 0.1 seconds) and lacks a clearly identifiable inspiral phase usually observed in the mergers of BBHs. This lack raises questions about its physical origin [8]. Though the LVK collaboration has interpreted GW190521 as the merger of BBHs with masses in $85^{+21}_{-14}M_{\odot}$ and $66^{+17}_{-18}M_{\odot}$, resulting in a postmerger remnant of $142^{+28}_{-16}M_{\odot}$ [7], see also relevant e.g. Refs. [9, 10, 11, 12], it is in tension with established astrophysical models of stellar evolution. Thus some alternative interpretations such as primordial black holes, cosmic strings, new light particles, and even horizonless compact objects have been presented [13, 14, 15, 16, 18, 17].
@@
In parallel, to investigate possible macroscopic signatures of quantum gravity and the longstanding black hole information paradox, a variety of horizonless exotic compact objects have been investigated, e.g.[19] for a recent review. The wormhole represents such an object connecting either two separate universes or two distant regions in a single universe through a throat [20, 21], see also [22, 23, 24, 25, 26, 27]. It has been shown that if the remnant object after BBHs merged is a wormhole, the corresponding ringdown burst propagating toward the other side of wormhole can be reflected back by the photon sphere on the far end, yielding a sequence of delayed pulses, known as echoes [28]. The properties of echoes, such as their amplitude and time delay, carry information about the underlying physical characteristics of corresponding wormholes, e.g.[29, 30, 31, 32, 33, 34, 35, 36, 37] ; see also Refs. [38, 39, 40, 41, 42].
@@
2 Model
2.1 Our echo-for-wormhole model
@@
As a proof-of-principle model, we consider a Schwarzschild–like Morris–Thorne wormhole [20, 21], see also recent [28], in which the Schwarzschild metrics of both sides are glued at $r=r_{0}>r_{S}$, and $r_{0}$ is the radius of wormhole throat and $r_{S}$ is the Schwarzschild radius. We note that, the remnant of GW190521 is highly spinning with $\chi_{f}=0.72^{+0.09}_{-0.12}$ [7]. Here, we adopt a nonspinning setup for simplicity and thus do not model spin-induced modifications to the effective potential or the echo spectrum, which constitutes a limitation of our model. As shown exactly in Ref.[30], the relevant issue on post-ringdown GW signals can be regarded equivalently as a one-dimensional scattering problem with a frequency-domain wave equation:
$\frac{d^{2}{\tilde{\Psi}}}{d{r_{*}}^{2}}+\left(\omega^{2}+V(r_{*},\omega)\right){\tilde{\Psi}}=0$ (2.1)
@@
In our model, see Fig.1, the BBHs inspiraled and merged in the left-side universe, and their postmerger object is a Schwarzschild-like wormhole connecting it to our universe. As a result, the postmerger GW ringdown signal, which is dominated by the frequency content of light-ring perturbations, will not only propagate outward at the speed of light, but also enter the wormhole throat and penetrate the right-side barrier into our universe. It can be expected that the reflections of such a pair of mirror barriers for the ringdown pulse will yield a series of echo pulses with sequentially suppressed amplitudes, and the frequency of each echo decreases until it reaches the wormhole quasinormal modes (QNMs) [19]. Therefore, the echo signal can be described as a sum of wormhole QNMs at late times [30]
$\Psi(t)\approx e^{-i\omega_{0}t}\left(\sum_{n=-\infty}^{+\infty}c_{n}e^{-n\pi t/L}e^{{\rm Im}(\omega_{n})t}\right),$ (2.2)
@@
$L\simeq 2\int_{r_{0}}^{\frac{3r_{S}}{2}}\frac{dr}{1-r_{S}/r}$ (2.3)
represents the separation between the mirror barriers, $\omega_{n}$ is the QNF of wormhole, and the coefficients $c_{n}$ can be written as
$c_{n}={\frac{1}{2L}}\int_{0}^{2L}dt\Psi_{1^{\rm st}{\rm echo}}e^{i\omega_{n}t},$ (2.4)
where $\Psi_{1^{\rm st}{\rm echo}}$ is the waveform of the first echo. The interval between two echoes is approximately $\Delta t_{\pm echo}\simeq 2L$, however, if the postmerger object is a wormhole collapsing into a black hole, the change of interval after a period is [31]
$\Delta t_{\rm echo}\simeq{2L\over 1-{\Delta L\over L}},$ (2.5)
@@
The postmerger ringdown signal closely resembles that of a black hole with Schwarzschild radius $r_{S}$. The waveform of the first echo is primarily governed by the wormhole QNF that lies closest to the corresponding black hole QNF $\omega_{0}^{\rm BH}$, thus it is natural to model the first echo as a Gaussian-like wave packet
$\Psi_{1^{\rm st}{\rm echo}}\sim e^{i\omega_{0}^{\rm BH}(t-t_{*})}e^{-{(t-t_{*})^{2}\over 2\beta^{2}}},$ (2.6)
where $\beta$ characterizes the width of the Gaussian-shaped echo pulse [43]. This result has been also verified numerically in Ref. [30].
2.2 Template of waveform
It can be postulated that, after the merger of binary black holes, a postmerger wormhole forms as an intermediate state, which rapidly pinches off (after the postmerger ringdown signal penetrates the right-side barrier into our universe) and eventually collapses into a black hole [31], or that the amplitudes of subsequent echoes fall below the sensitivity threshold of current detectors [18]. In either case, only the first echo pulse can be detected in our universe [^1] Recently, numerous studies have focused on the search for echoes after the BBHs inspiraled and merged in one single universe, e.g.[59, 60, 70, 61, 62, 63, 64, 65, 66, 18, 67, 68, 69], showing no evidence for echo, however, here what we consider is to take the GW190521 event itself as an echo (detected in our universe) of postmerger wormhole in another universe in which the BBHs inspiraled and merged.. According to (2.6), we use a simplified sine-Gaussian template [^2] The sine-Gaussian template has been widely used in the search for GW echoes, e.g.Refs.[71, 72] and LIGO-Virgo-KAGRA collaboration’s work with GWTC-3 [73].:
$h(t)$ $\left(\sim e^{2\pi if_{c}(t-t_{c})}e^{-{(t-t_{c})^{2}\over 2\beta^{2}}}\right)$ (2.7)
$=$ $\mathcal{A}\mathrm{cos}\left(2\pi f_{c}(t-t_{c})+\varphi\right)e^{-\frac{(t-t_{c})^{2}}{2\beta^{2}}},$
@@ 2.2 Template of waveform
The amplitude $\mathcal{A}$ depends on the energy of the GWs emitted by the BBHs merger in another universe. However, in our case, the total energy emitted by the source is unknown, which causes the amplitude parameter inferred from the Bayesian analysis to encode both the intrinsic strength of the source and its distance to the detector. To facilitate the parameter estimation, we fix a reference strain amplitude $A_{ref}=10^{-19}$ at a fiducial distance of 1Mpc, the physical amplitude $\mathcal{A}$ is then rescaled by a dimensionless factor $A_{m}=\mathcal{A}/A_{ref}$ and set as a free parameter in our analysis. This normalization not only improves numerical stability but also yields a consistent scale for comparing with LVK BBHs models.
3 Results
3.1 Our echo-for-wormhole model
@@
Table 1: Prior ranges of parameters for our echo-for-wormhole waveform (2.7). Here, $\psi$ is the polarization, while $\iota$ is the inclination refers to the orientation of binary relative to the line of sight. Geocent time parameter specifies the arrival time of the echo pulse at the Earth. The physical interpretations of other parameters are presented in Sec.2.2
@@
The priors of our echo-for-wormhole model are shown in Tab.1. To avoid an implicit prior preference for large amplitudes, we adopt a log-uniform prior for $A_{m}$ within the specified range, while all other parameters are assigned uniform priors within their respective ranges. In the initial parameter estimation, we find that the maximum-posterior sky location inferred is consistent with that obtained using the LVK BBHs model, and also with the localization reported for ZTF19abanrhr, as shown in Appendix A. Thus we fix the location parameters with ZFT19abanrhr localization (RA, DEC)=(3.35847,0.60781)=(3.35847,~0.60781) to break the degeneracy between arrival time and sky position. The marginalized posterior probability distributions for all parameters in our model are presented in Fig. 2, where $f_{c}=56.93^{+1.63}_{-1.86}$ Hz and $\beta\approx 0.02$. This result is consistent with the narrow-band feature of an echo pulse produced by a wormhole in the frequency domain.
In Fig.3, we plot the best-fit echo-for-wormhole waveform (2.7) with the strain data in L1 detector in time-domain, and also the BBHs waveform for a comparison, where owing to the short duration of GW190521 (approximately $0.1$ s), we additionally bandpass the strain data between 20–256 Hz to suppress high frequencies, resulting in an almost 5-cycle waveform. Here, to facilitate a direct comparison between the two models, we whiten both the strain data and the corresponding waveforms according to
$\tilde{h}_{w}(f)=\frac{\tilde{h}(f)}{\sqrt{\tilde{S}_{n}(f)}},$ (3.1)
where $\tilde{h}_{w}(f)$ denotes the whitened waveform in the frequency domain; $\tilde{S}_{n}(f)$ is the one-sided noise power spectral density (PSD) estimated from the Gravitational Wave Open Science Center (GWOSC) strain data [44] using the Welch method implemented in PyCBC with 4 s segments [45].final
@@
3.2 Bayesian comparison with BBHs model
Figure 4: The posterior distribution of intrinsic parameters and luminosity distance for the standard BBHs IMR event with IMRPhenomXPHM waveform. This distribution is consistent with the result reported by the LVK collaboration for GW190521 using the same waveform model [11].
Here, we use the IMRPhenomXPHM as the BBHs waveform template [^3] Though the waveform models NRSur7dq4 [74] and SEOBNRv4PHM [75, 76, 77, 78] might yield more precise parameter estimates, the differences between different BBHs models are negligible compared to the impact of distinct source hypotheses on the Bayesian analysis., as it accommodates high mass ratios and spin precession, effectively covering the parameter space relevant to GW190521 [46]. In addition, its frequency-domain construction offers high computational efficiency, facilitating a systematic comparison with our echo-for-wormhole model. We adopt the same prior distributions as those used in Ref. [11], this choice places greater emphasis on unequal–mass systems, thereby improving sampling convergence in the unequal–mass regime. And using the IMRPhenomXPHM waveform we recover posterior of relevant parameters [^4] Here, we also used the localization reported for ZTF19abanrhr, see Appendix A. However, we have verified that the results are actually nearly same with those without the ZTF19abanrhr localization. (see Fig.4), which consistent with the parameter estimates reported by the LVK collaboration [11], the merger of BBHs with masses of $m_{1}=97^{+32}_{-21}\mathrm{M}_{\odot}$ and $m_{2}=59^{+22}_{-25}\mathrm{M}_{\odot}$ at a luminosity distance $d_{L}=3.5^{+2.4}_{-2.0}$Gpc.
To evaluate the plausibility of our model as a potential explanation for the source of GW190521, we compute the SNRs of the BBHs model and our model for each detector, as well as the network SNR, by using the publicly available PyCBC software [45]. The results are summarized in Table 2. The network SNR is defined as the quadrature sum of the individual SNRs across the detectors,
$\text{SNR}^{2}_{\text{net}}=\sum_{ifo=\text{H1,L1,V1}}\text{SNR}^{2}_{ifo},$ (3.2)